OpenStudy (anonymous):

TRIG PROBLEM! i give medals. prove the identity is true:

OpenStudy (anonymous):

\[\sin ^{2}x (\cot ^{2}x+1) = 1\]

OpenStudy (anonymous):

Since cot(x) = 1/tan(x) and tan(x) = sin(x)/cos(x) then sin^2(x)*(1 + cot^2(x)) = sin^2(x)*(1 + cos^2(x)/sin^2(x)) = sin^2(x) + cos^2(x) = 1 since sin^2(x) + cos^2(x) =1

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

okay @sophiaedge , that makes sense, what next?

OpenStudy (anonymous):

Mmmm I don't know I emailed that to my friend and that's all he sent back sorry!

OpenStudy (anonymous):

aw okay thanks

OpenStudy (anonymous):

can anyone explain this to me? :(

OpenStudy (jhannybean):

\[\sin ^{2}x (\cot ^{2}x+1) = 1\]\[\sin^2(x)\left(\frac{\cos^2(x)}{\sin^2(x)} +1\right) = 1\]\[\sin^2\cdot \frac{\cos^2(x)}{\sin^2(x)} + \sin^2(x) = 1\]\[\cancel{\sin^2(x)}\cdot \frac{\cos^2(x)}{\cancel{\sin^2(x)}} + \sin^2(x) = 1\]\[\cos^2(x) + \sin^2(x)=1\]\[\boxed{1 = 1}~~\checkmark\]

OpenStudy (anonymous):

oh my god. i love you.

OpenStudy (anonymous):

i have one more, i will post in a different question, is that okay?

OpenStudy (anonymous):

that way i can give you a medal.. id give you a trophy if i could, but i can only do medals

OpenStudy (jhannybean):

Haha, sure. although @sophiaedge is the one who originally solved it :P